OpenStudy (anonymous):

if p(z)=z^3-az^2+bz-4 find the values of (a) and (b) if p(2)=0

OpenStudy (anonymous):

hope this helps p(z)=z^3+az^2+bz+c p(-2)=-8+4a-2b+c=0=> 4a-2b+c=8--------------(1) If z is a root, then z divides p(z) =>the remainder (9+5a-3b+c)+(-12a+2b+46)i=0 =>5a-3b+c=-9----------(2) .-12a+2b=-46 =>6a-b=23-------------(3) Solving the system of (1),(2) & (3) for a,b,c get a=8 b=25 c=26

OpenStudy (anonymous):

did you work that out ?

OpenStudy (anonymous):

yes